# FlashRNN: Optimizing traditional RNNs on modern hardware

### v2

@@ 1 Introduction
Sequence models are at the core of many applications like time-series modeling, natural language processing, text, audio and video models, and predictions for physical systems based on ODEs or PDEs (Vaswani et al., 2017; Degrave et al., 2022; Nearing et al., 2024). While there are modern sequence-parallelizable architectures like the Transformer (Vaswani et al., 2017), Mamba (Gu and Dao, 2023) or mLSTM (Beck et al., 2024), these lack the state-tracking capabilities (Merrill et al., 2024; Merrill and Sabharwal, 2023; Delétang et al., 2023) of traditional RNNs like LSTM (Hochreiter and Schmidhuber, 1997), GRU (Cho et al., 2014) and sLSTM (Beck et al., 2024).
@@
Our library facilitates research in the direction of state-tracking enabled RNN architectures, in two ways: Firstly, it enables easier and more efficient use of recent RNN-architectures like sLSTM (Beck et al., 2024). This includes the notion of block-diagonal recurrent matrices that can speed up networks while lowering the number of parameters. Secondly, it can be easily extended to novel RNN-like architecture variants, as it supports generic state and gate numbers per cell. The LSTM (Hochreiter and Schmidhuber, 1997; F.A. Gers, 1999), with its two states and four gates (we consider the cell update as a fourth "gate" for simplicity here), can be implemented as easy as a simple Elman-RNN with one state and one gate (Elman, 1990), or sLSTM with its three states and four gates (Beck et al., 2024).
@@ 2 Related work
Recently, novel sequence models taking inspiration of Linear Attention (Katharopoulos et al., 2020) have shown promising performance compared to Transformer Attention (Beck et al., 2024; Yang et al., 2024; Dao and Gu, 2024). Yang et al. (2024) provide an hardware-efficient algorithm and implementation in Triton for Gated Linear Attention that trades off memory movement against parallelizability and show that it is faster than FlashAttention2. Traditional RNNs like LSTMs (Hochreiter and Schmidhuber, 1997) or GRUs (Cho et al., 2014) are still widely used in many applications, such as for example time series modeling or reinforcement learning (Nearing et al., 2024; Degrave et al., 2022). Many of these applications rely on optimized closed-source implementations of these RNN operations such as in the NVIDIA cuDNN  [^1] https://developer.nvidia.com/cudnn library, which is integrated in PyTorch. Sharvil (2020) provide an open-source alternative in CUDA for specific LSTM and GRU variants in their HASTE library, which served as inspiration for this work. HASTE is limited in speed due to a sequence of alternating calls of matrix multiplication and point-wise kernels, as well as its limitation to higher (but slower) precision.
@@ 3 Generic Recurrent Neural Network architecture with memory mixing
A generic RNN architecture that we aim to optimize has $N_{s}$ states ${\bm{s}}^{(i)}\in\mathbb{R}^{d}$, and $N_{g}$ gates (or pre-activations) ${\bm{g}}^{(j)}\in\mathbb{R}^{d}$, with $d$ being the embedding dimension or hidden size of the RNN. For example the LSTM (Hochreiter and Schmidhuber, 1997) has $N_{s}=2$ states and $N_{g}=4$ gates.
@@
${\bm{s}}^{(i)}_{t}$ $={\mathcal{P}}^{(i)}\left(\left({\bm{s}}^{(i^{\prime})}_{t-1}\right)_{i^{\prime}},\left({\bm{g}}^{(j)}_{t}\right)_{j}\right),$ (2)
@@ 4 Generic gradient for back-propagation through time
$\delta{\bm{g}}^{(j)}_{t}$ $=\frac{\partial{\mathcal{P}}^{(l)}\left(\left({\bm{s}}^{(k)}_{t-1}\right)_{k},\left({\bm{g}}^{(j)}_{t}\right)_{j}\right)}{\partial{\bm{g}}^{(j)}_{t}}\delta{\bm{s}}^{(l)}_{t}$ (3)
$\delta{\bm{s}}^{(i)}_{t-1}$ $=\frac{\partial{\mathcal{P}}^{(l)}\left(\left({\bm{s}}^{(k)}_{t-1}\right)_{k},\left({\bm{g}}^{(j)}_{t}\right)_{j}\right)}{\partial{\bm{s}}^{(i)}_{t-1}}\delta{\bm{s}}^{(l)}_{t}+\left({{\bm{R}}^{(j)}}^{T}\delta{\bm{g}}^{(j)}_{t-1}\qquad\text{if}\;\;i=0\right)$ (4)
@@ 4.1 Vanishing and Exploding gradients and Gradient Modifications
For a neural network to be stably trainable, there must not be exploding gradients, also vanishing gradients should be prohibited for long context sequence modeling (Hochreiter and Schmidhuber, 1997). Still, for the generic structure of Equations 3, there can be exploding components: Firstly, one or more eigenvalues of the point-wise function Jacobian can be greater than one in magnitude. This can be mitigated by a proper choice of the point-wise function. Secondly, the combination of recurrent matrix and gate gradients with the gradient $\frac{\partial{\mathcal{P}}^{(0)}}{\partial{\bm{g}}^{(j)}_{t}}{{\bm{R}}^{(j)}}^{T}$ could have singular values of magnitude $>1$. This case cannot be excluded directly, as the recurrent matrix consists of trainable weights with usually unconstrained magnitude. However, for practical training this is rarely a limitation.
@@ 4.2 Head-wise parallelization
Many more recent architectures also rely on this head-wise parallelization primitive (Beck et al., 2024; Yang et al., 2024; Dao and Gu, 2024), where the embedding or hidden vector of dimension $d$ is split into $N_{heads}$ heads of smaller dimension $d_{head}=d/N_{head}$, each of which is processed independently inside the sequential part. In FlashRNN, we apply this primitive to traditional RNNs by dividing the recurrent matrix ${\bm{R}}$ into multiple blocks or heads ${\bm{R}}_{head}\in\mathbb{R}^{d_{head}\times d_{head}}$ rendering the recurrent matrix ${\bm{R}}$ as a block-diagonal matrix.
@@ 6.3 State Tracking Task
To show state tracking capabilities of traditional RNNs in contrast to Transformers and State Space Models experimentally, we train our implementation on the Parity task and evaluate on longer sequences to measure extrapolation capabilities (Zhou et al., 2024). This serves as a litmus test for state tracking capabilities (Merrill et al., 2024).
@@ Acknowledgments
We thank Markus Spanring, Maxim Milakov, Pieter-Jan Hoedt, Günter Klambauer and Fabian Paischer for helpful discussions and feedback.
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, Leonardo at CINECA, Italy and LUMI at CSC, Finland. The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects Medical Cognitive Computing Cen- ter (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF- 36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT- 2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Compe- tence Center Hagenberg GmbH, Borealis AG, T ¨UV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
@@ Appendix A RNN variants with memory mixing / recurrent connections modeled in FlashRNN
[Hochreiter and Schmidhuber, 1997, F.A. Gers, 1999] Number of states: 2, Number of gates: 4

### v3

# FlashRNN: I/O-Aware Optimization of Traditional RNNs on modern hardware

@@ 1 Introduction
Sequence models are at the core of many applications like time-series modeling, natural language processing, text, audio and video models, and predictions for physical systems based on ODEs or PDEs (Vaswani et al., 2017; Degrave et al., 2022; Nearing et al., 2024). While there are modern sequence-parallelizable architectures like the Transformer (Vaswani et al., 2017), Mamba (Gu & Dao, 2023) or mLSTM (Beck et al., 2024), these lack the state-tracking capabilities (Merrill et al., 2024; Merrill & Sabharwal, 2023; Delétang et al., 2023) of traditional RNNs like LSTM (Hochreiter & Schmidhuber, 1997), GRU (Cho et al., 2014) and sLSTM (Beck et al., 2024).
@@
Our library facilitates research in the direction of state-tracking enabled RNN architectures, in two ways: Firstly, it enables easier and more efficient use of recent RNN-architectures like sLSTM (Beck et al., 2024). This includes the notion of block-diagonal recurrent matrices that can speed up networks while lowering the number of parameters. Secondly, it can be easily extended to novel RNN-like architecture variants, as it supports generic state and gate numbers per cell. The LSTM (Hochreiter & Schmidhuber, 1997; Gers et al., 1999), with its two states and four gates (we consider the cell update as a fourth "gate" for simplicity here), can be implemented as easy as a simple Elman-RNN with one state and one gate (Elman, 1990), or sLSTM with its three states and four gates (Beck et al., 2024).
@@ 2 Related work
Recently, novel sequence models taking inspiration of Linear Attention (Katharopoulos et al., 2020) have shown promising performance compared to Transformer Attention (Beck et al., 2024; Yang et al., 2024; Dao & Gu, 2024). Yang et al. (2024) provide an hardware-efficient algorithm and implementation in Triton for Gated Linear Attention that trades off memory movement against parallelizability and show that it is faster than FlashAttention2.
Traditional RNNs like LSTMs (Hochreiter & Schmidhuber, 1997) or GRUs (Cho et al., 2014) are still widely used in many applications, such as for example time series modeling or reinforcement learning (Nearing et al., 2024; Degrave et al., 2022). Many of these applications rely on optimized closed-source implementations of these RNN operations such as in the NVIDIA cuDNN  [^1] https://developer.nvidia.com/cudnn library, which is integrated in PyTorch. Sharvil (2020) provide an open-source alternative in CUDA for specific LSTM and GRU variants in their HASTE library, which served as inspiration for this work. HASTE is limited in speed due to a sequence of alternating calls of matrix multiplication and point-wise kernels, as well as its limitation to higher (but slower) precision.
@@ 3 Generic Recurrent Neural Network architecture with memory mixing
A generic RNN architecture that we aim to optimize has $N_{s}$ states ${\bm{s}}^{(i)}\in\mathbb{R}^{d}$, and $N_{g}$ gates (or pre-activations) ${\bm{g}}^{(j)}\in\mathbb{R}^{d}$, with $d$ being the embedding dimension or hidden size of the RNN. For example the LSTM (Hochreiter & Schmidhuber, 1997) has $N_{s}=2$ states and $N_{g}=4$ gates.
@@
${\bm{s}}^{(i)}_{t}$ $={\mathcal{P}}^{(i)}\left(\left({\bm{s}}^{(i^{\prime})}_{t-1}\right)_{i^{\prime}\in\{1..N_{s}\}},\left({\bm{g}}^{(j)}_{t}\right)_{j\in\{1..N_{g}\}}\right),$ (2)
@@ 4 Generic gradient for back-propagation through time
$\delta{\bm{g}}^{(j)}_{t}$ $=\frac{\partial{\mathcal{P}}^{(l)}\left(\left({\bm{s}}^{(k)}_{t-1}\right)_{k\in\{1..N_{s}\}},\left({\bm{g}}^{(j)}_{t}\right)_{j\in\{1..N_{g}\}}\right)}{\partial{\bm{g}}^{(j)}_{t}}\delta{\bm{s}}^{(l)}_{t}$ (3)
$\delta{\bm{s}}^{(i)}_{t-1}$ $=\frac{\partial{\mathcal{P}}^{(l)}\left(\left({\bm{s}}^{(k)}_{t-1}\right)_{k\in\{1..N_{s}\}},\left({\bm{g}}^{(j)}_{t}\right)_{j\in\{1..N_{g}\}}\right)}{\partial{\bm{s}}^{(i)}_{t-1}}\delta{\bm{s}}^{(l)}_{t}+\left(\sum_{j\in\{1..N_{g}\}}{{\bm{R}}^{(j)}}^{T}\delta{\bm{g}}^{(j)}_{t-1}\qquad\text{if}\;\;i=0\right)$ (4)
@@ 4.1 Vanishing and Exploding gradients and Gradient Modifications
For a neural network to be stably trainable, there must not be exploding gradients, also vanishing gradients should be prohibited for long context sequence modeling (Hochreiter & Schmidhuber, 1997). Still, for the generic structure of Equations 3, there can be exploding components: Firstly, one or more eigenvalues of the point-wise function Jacobian can be greater than one in magnitude. This can be mitigated by a proper choice of the point-wise function. Secondly, the combination of recurrent matrix and gate gradients with the gradient $\frac{\partial{\mathcal{P}}^{(0)}}{\partial{\bm{g}}^{(j)}_{t}}{{\bm{R}}^{(j)}}^{T}$ could have singular values of magnitude $>1$. This case cannot be excluded directly, as the recurrent matrix consists of trainable weights with usually unconstrained magnitude. However, for practical training this is rarely a limitation.
@@ 4.2 Head-wise parallelization
Many more recent architectures also rely on this head-wise parallelization primitive (Beck et al., 2024; Yang et al., 2024; Dao & Gu, 2024), where the embedding or hidden vector of dimension $d$ is split into $N_{heads}$ heads of smaller dimension $d_{head}=d/N_{head}$, each of which is processed independently inside the sequential part. In FlashRNN, we apply this primitive to traditional RNNs by dividing the recurrent matrix ${\bm{R}}$ into multiple blocks or heads ${\bm{R}}_{head}\in\mathbb{R}^{d_{head}\times d_{head}}$ rendering the recurrent matrix ${\bm{R}}$ as a block-diagonal matrix.
@@ 6.3 State Tracking Task
To show state tracking capabilities of traditional RNNs in contrast to Transformers and State Space Models experimentally, we train our implementation on the Parity task and evaluate on longer sequences to measure extrapolation capabilities (Zhou et al., 2024). This serves as a litmus test for the simplest subclass of state tracking capabilities (Merrill et al., 2024).
@@ Acknowledgments
We thank Markus Spanring, Maxim Milakov (NVIDIA), Pieter-Jan Hoedt, Günter Klambauer and Fabian Paischer for helpful discussions and feedback.
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, Leonardo at CINECA, Italy and LUMI at CSC, Finland. The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/COE12]. We thank the projects INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank NXAI GmbH, Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
@@ Appendix A RNN variants with memory mixing / recurrent connections modeled in FlashRNN
[Hochreiter & Schmidhuber, 1997, Gers et al., 1999] Number of states: 2, Number of gates: 4